### v2

@@ Appendix A Extended Data
| ResNet50 | 0.700±0.165 | 0.683±0.178 | 0.888±0.070 |
| Phikon | 0.804±0.158 | 0.801±0.176 | 0.941±0.057 |
| Ctranspath | 0.789±0.119 | 0.791±0.124 | 0.931±0.060 |
| UNI | 0.814±0.148 | 0.815±0.155 | 0.941±0.065 |
| CONCH | 0.803±0.152 | 0.800±0.167 | 0.942±0.052 |
| PLIP | 0.772±0.141 | 0.768±0.142 | 0.915±0.055 |
| GPFM | 0.833±0.142 | 0.834±0.150 | 0.956±0.046 |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.845 (0.800-0.893) | 0.842 (0.791-0.893) | 0.929 (0.896-0.959) |
| Phikon | 0.888 (0.845-0.928) | 0.885 (0.843-0.924) | 0.982 (0.968-0.993) |
| Ctranspath | 0.894 (0.850-0.934) | 0.895 (0.851-0.933) | 0.963 (0.936-0.985) |
| UNI | 0.928 (0.891-0.961) | 0.928 (0.890-0.962) | 0.977 (0.957-0.992) |
| CONCH | 0.924 (0.888-0.957) | 0.924 (0.881-0.957) | 0.986 (0.971-0.996) |
| PLIP | 0.865 (0.821-0.908) | 0.865 (0.812-0.909) | 0.942 (0.910-0.969) |
| GPFM | 0.948 (0.915-0.976) | 0.947 (0.918-0.976) | 0.986 (0.973-0.996) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.847 (0.803-0.871) | 0.847 (0.803-0.871) | 0.937 (0.899-0.945) |
| Phikon | 0.901 (0.873-0.928) | 0.900 (0.880-0.925) | 0.967 (0.951-0.980) |
| Ctranspath | 0.887 (0.858-0.916) | 0.887 (0.856-0.914) | 0.965 (0.950-0.977) |
| UNI | 0.911 (0.883-0.937) | 0.911 (0.884-0.939) | 0.960 (0.942-0.976) |
| CONCH | 0.876 (0.844-0.903) | 0.876 (0.844-0.905) | 0.961 (0.944-0.975) |
| PLIP | 0.841 (0.805-0.876) | 0.841 (0.808-0.873) | 0.939 (0.918-0.957) |
| GPFM | 0.906 (0.877-0.932) | 0.906 (0.880-0.934) | 0.974 (0.961-0.985) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.870 (0.805-0.926) | 0.858 (0.797-0.910) | 0.963 (0.937-0.983) |
| Phikon | 0.964 (0.936-0.987) | 0.960 (0.937-0.982) | 0.995 (0.989-0.999) |
| Ctranspath | 0.883 (0.820-0.938) | 0.888 (0.833-0.939) | 0.981 (0.965-0.992) |
| UNI | 0.903 (0.847-0.954) | 0.913 (0.862-0.957) | 0.982 (0.966-0.995) |
| CONCH | 0.941 (0.891-0.979) | 0.937 (0.896-0.974) | 0.988 (0.977-0.997) |
| PLIP | 0.899 (0.845-0.947) | 0.904 (0.853-0.948) | 0.980 (0.963-0.993) |
| GPFM | 0.925 (0.874-0.967) | 0.930 (0.885-0.966) | 0.996 (0.992-0.999) |
@@
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.658 (0.585-0.735) | 0.691 (0.596-0.783) | 0.846 (0.768-0.911) |
| Phikon | 0.794 (0.718-0.865) | 0.835 (0.751-0.901) | 0.936 (0.887-0.977) |
| Ctranspath | 0.843 (0.767-0.914) | 0.859 (0.790-0.917) | 0.931 (0.870-0.975) |
| UNI | 0.869 (0.797-0.929) | 0.879 (0.810-0.932) | 0.946 (0.894-0.987) |
| CONCH | 0.835 (0.750-0.905) | 0.875 (0.807-0.934) | 0.944 (0.902-0.979) |
| PLIP | 0.823 (0.747-0.897) | 0.820 (0.750-0.888) | 0.893 (0.824-0.950) |
| GPFM | 0.881 (0.813-0.947) | 0.907 (0.850-0.956) | 0.950 (0.898-0.990) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.568 (0.515-0.615) | 0.522 (0.463-0.571) | 0.835 (0.776-0.892) |
| Phikon | 0.707 (0.621-0.797) | 0.701 (0.602-0.800) | 0.898 (0.852-0.942) |
| Ctranspath | 0.674 (0.592-0.757) | 0.664 (0.559-0.754) | 0.908 (0.871-0.946) |
| UNI | 0.746 (0.660-0.840) | 0.738 (0.640-0.824) | 0.913 (0.865-0.956) |
| CONCH | 0.677 (0.606-0.752) | 0.668 (0.575-0.771) | 0.923 (0.883-0.958) |
| PLIP | 0.679 (0.596-0.773) | 0.676 (0.579-0.782) | 0.866 (0.805-0.917) |
| GPFM | 0.749 (0.660-0.834) | 0.758 (0.658-0.841) | 0.936 (0.896-0.965) |

### v3

@@ Appendix A Extended Data
| ResNet50 | 0.563±0.183 | 0.518±0.215 | 0.769±0.135 |
| Phikon | 0.693±0.191 | 0.670±0.228 | 0.868±0.105 |
| Ctranspath | 0.698±0.183 | 0.677±0.212 | 0.855±0.111 |
| UNI | 0.721±0.181 | 0.706±0.214 | 0.875±0.105 |
| CONCH | 0.695±0.180 | 0.664±0.214 | 0.849±0.140 |
| PLIP | 0.668±0.156 | 0.639±0.186 | 0.835±0.109 |
| CHIEF | 0.687±0.171 | 0.667±0.202 | 0.856±0.108 |
| Prov-Gigapath | 0.703±0.173 | 0.676±0.209 | 0.854±0.113 |
| GPFM | 0.752±0.161 | 0.736±0.179 | 0.891±0.096 |
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | TCGA-NSCLC | 0.845 (0.800-0.893) | 0.842 (0.791-0.893) | 0.929 (0.896-0.959) |
| Phikon | TCGA-NSCLC | 0.888 (0.845-0.928) | 0.885 (0.843-0.924) | 0.982 (0.968-0.993) |
| Ctranspath | TCGA-NSCLC | 0.894 (0.850-0.934) | 0.895 (0.851-0.933) | 0.963 (0.936-0.985) |
| UNI | TCGA-NSCLC | 0.928 (0.891-0.961) | 0.928 (0.890-0.962) | 0.977 (0.957-0.992) |
| CONCH | TCGA-NSCLC | 0.924 (0.888-0.957) | 0.924 (0.881-0.957) | 0.986 (0.971-0.996) |
| PLIP | TCGA-NSCLC | 0.865 (0.821-0.908) | 0.865 (0.812-0.909) | 0.942 (0.910-0.969) |
| CHIEF | TCGA-NSCLC | 0.910 (0.874-0.951) | 0.910 (0.866-0.943) | 0.971 (0.951-0.988) |
| Prov-Gigapath | TCGA-NSCLC | 0.918 (0.880-0.953) | 0.919 (0.877-0.957) | 0.967 (0.942-0.987) |
| GPFM | TCGA-NSCLC | 0.948 (0.915-0.976) | 0.947 (0.918-0.976) | 0.986 (0.973-0.996) |
| ResNet50 | CPTAC-NSCLC | 0.847 (0.803-0.871) | 0.847 (0.803-0.871) | 0.937 (0.899-0.945) |
| Phikon | CPTAC-NSCLC | 0.901 (0.873-0.928) | 0.900 (0.880-0.925) | 0.967 (0.951-0.980) |
| Ctranspath | CPTAC-NSCLC | 0.887 (0.858-0.916) | 0.887 (0.856-0.914) | 0.965 (0.950-0.977) |
| UNI | CPTAC-NSCLC | 0.911 (0.883-0.937) | 0.911 (0.884-0.939) | 0.960 (0.942-0.976) |
| CONCH | CPTAC-NSCLC | 0.876 (0.844-0.903) | 0.876 (0.844-0.905) | 0.961 (0.944-0.975) |
| PLIP | CPTAC-NSCLC | 0.841 (0.805-0.876) | 0.841 (0.808-0.873) | 0.939 (0.918-0.957) |
| CHIEF | CPTAC-NSCLC | 0.881 (0.851-0.909) | 0.882 (0.851-0.912) | 0.964 (0.949-0.978) |
| Prov-Gigapath | CPTAC-NSCLC | 0.882 (0.853-0.911) | 0.883 (0.853-0.911) | 0.971 (0. 957-0.982) |
| GPFM | CPTAC-NSCLC | 0.906 (0.877-0.932) | 0.906 (0.880-0.934) | 0.974 (0.961-0.985) |
| ResNet50 | Center-1-NSCLC* | 0.492 (0.481-0.5) | 0.457 (0.443-0.471) | 0.611 (0.497-0.726) |
| Phikon | Center-1-NSCLC* | 0.566 (0.497-0.644) | 0.590 (0.475-0.713) | 0.660 (0.526-0.770) |
| Ctranspath | Center-1-NSCLC* | 0.695 (0.583-0.802) | 0.695 (0.590-0.789) | 0.722 (0.592-0.848) |
| UNI | Center-1-NSCLC* | 0.764 (0.677-0.857) | 0.788 (0.703-0.864) | 0.819 (0.729-0.909) |
| CONCH | Center-1-NSCLC* | 0.617 (0.512-0.724) | 0.579 (0.502-0.658) | 0.595 (0.439-0.749) |
| PLIP | Center-1-NSCLC* | 0.682 (0.577-0.783) | 0.602 (0.526-0.684) | 0.770 (0.667-0.864) |
| CHIEF | Center-1-NSCLC* | 0.693 (0.575-0.814) | 0.571 (0.496-0.642) | 0.746 (0.592-0.878) |
| Prov-Gigapath | Center-1-NSCLC* | 0.725 (0.598-0.864) | 0.759 (0.624-0.852) | 0.727 (0.554-0.900) |
| GPFM | Center-1-NSCLC* | 0.614 (0.534-0.689) | 0.653 (0.541-0.750) | 0.823 (0.728-0.902) |
| cls | Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 2 | Center-1-LMD | 0.816 (0.752-0.876) | 0.819 (0.759-0.880) | 0.905 (0.850-0.950) |
| Phikon | 2 | Center-1-LMD | 0.902 (0.853-0.945) | 0.894 (0.844-0.943) | 0.984 (0.967-0.996) |
| Ctranspath | 2 | Center-1-LMD | 0.913 (0.863-0.956) | 0.908 (0.858-0.957) | 0.981 (0.962-0.994) |
| UNI | 2 | Center-1-LMD | 0.908 (0.860-0.952) | 0.901 (0.852-0.950) | 0.981 (0.962-0.997) |
| CONCH | 2 | Center-1-LMD | 0.910 (0.861-0.958) | 0.908 (0.852-0.951) | 0.965 (0.935-0.990) |
| PLIP | 2 | Center-1-LMD | 0.838 (0.777-0.896) | 0.837 (0.768-0.894) | 0.940 (0.901-0.972) |
| CHIEF | 2 | Center-1-LMD | 0.930 (0.882-0.972) | 0.935 (0.890-0.971) | 0.983 (0.965-0.996) |
| Prov-Gigapath | 2 | Center-1-LMD | 0.833 (0.770-0.894) | 0.841 (0.772-0.897) | 0.968 (0.936-0.992) |
| GPFM | 2 | Center-1-LMD | 0.940 (0.902-0.977) | 0.943 (0.904-0.978) | 0.985 (0.966-0.998) |
| ResNet50 | 2 | Center-2-LMD* | 0.588 (0.544-0.633) | 0.559 (0.509-0.608) | 0.651 (0.602-0.703) |
| Phikon | 2 | Center-2-LMD* | 0.801 (0.762-0.839) | 0.805 (0.763-0.839) | 0.914 (0.886-0.937) |
| Ctranspath | 2 | Center-2-LMD* | 0.800 (0.768-0.835) | 0.782 (0.740-0.821) | 0.885 (0.854-0.916) |
| UNI | 2 | Center-2-LMD* | 0.820 (0.783-0.855) | 0.819 (0.783-0.857) | 0.918 (0.893-0.939) |
| CONCH | 2 | Center-2-LMD* | 0.859 (0.828-0.888) | 0.849 (0.813-0.881) | 0.950 (0.931-0.967) |
| PLIP | 2 | Center-2-LMD* | 0.690 (0.645-0.733) | 0.691 (0.647-0.738) | 0.764 (0.721-0.807) |
| CHIEF | 2 | Center-2-LMD* | 0.748 (0.703-0.788) | 0.750 (0.708-0.794) | 0.881 (0.848-0.911) |
| Prov-Gigapath | 2 | Center-2-LMD* | 0.666 (0.624-0.708) | 0.661 (0.618-0.704) | 0.744 (0.698-0.787) |
| GPFM | 2 | Center-2-LMD* | 0.800 (0.763-0.838) | 0.805 (0.763-0.845) | 0.927 (0.903-0.948) |
| ResNet50 | 6 | Center-1-LMD | 0.378 (0.305-0.475) | 0.365 (0.283-0.453) | 0.895 (0.750-0.933) |
| Phikon | 6 | Center-1-LMD | 0.537 (0.433-0.666) | 0.539 (0.409-0.646) | 0.955 (0.806-0.979) |
| Ctranspath | 6 | Center-1-LMD | 0.640 (0.481-0.796) | 0.666 (0.516-0.788) | 0.959 (0.806-0.979) |
| UNI | 6 | Center-1-LMD | 0.709 (0.570-0.856) | 0.702 (0.564-0.821) | 0.961 (0.816-0.989) |
| CONCH | 6 | Center-1-LMD | 0.526 (0.472-0.637) | 0.475 (0.398-0.569) | 0.955 (0.811-0.985) |
| PLIP | 6 | Center-1-LMD | 0.600 (0.520-0.716) | 0.534 (0.436-0.638) | 0.936 (0.799-0.965) |
| CHIEF | 6 | Center-1-LMD | 0.640 (0.486-0.808) | 0.679 (0.505-0.829) | 0.959 (0.810-0.987) |
| Prov-Gigapath | 6 | Center-1-LMD | 0.702 (0.568-0.850) | 0.695 (0.530-0.819) | 0.967 (0.815-0.989) |
| GPFM | 6 | Center-1-LMD | 0.771 (0.640-0.924) | 0.767 (0.598-0.894) | 0.974 (0.821-0.991) |
| ResNet50 | 6 | Center-2-LMD* | 0.277 (0.234-0.330) | 0.250 (0.212-0.291) | 0.721 (0.674-0.765) |
| Phikon | 6 | Center-2-LMD* | 0.459 (0.394-0.524) | 0.429 (0.380-0.479) | 0.871 (0.831-0.904) |
| Ctranspath | 6 | Center-2-LMD* | 0.509 (0.441-0.569) | 0.519 (0.450-0.577) | 0.886 (0.851-0.917) |
| UNI | 6 | Center-2-LMD* | 0.568 (0.492-0.656) | 0.542 (0.487-0.592) | 0.900 (0.865-0.928) |
| CONCH | 6 | Center-2-LMD* | 0.525 (0.471-0.574) | 0.428 (0.388-0.471) | 0.906 (0.868-0.940) |
| PLIP | 6 | Center-2-LMD* | 0.405 (0.346-0.468) | 0.399 (0.347-0.445) | 0.861 (0.816-0.896) |
| CHIEF | 6 | Center-2-LMD* | 0.490 (0.419-0.557) | 0.532 (0.454-0.595) | 0.885 (0.854-0.916) |
| Prov-Gigapath | 6 | Center-2-LMD* | 0.551 (0.496-0.602) | 0.481 (0.429-0.525) | 0.853 (0.809-0.888) |
| GPFM | 6 | Center-2-LMD* | 0.591 (0.526-0.653) | 0.601 (0.541-0.655) | 0.908 (0.881-0.932) |
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | TCGA-RCC | 0.870 (0.805-0.926) | 0.858 (0.797-0.910) | 0.963 (0.937-0.983) |
| Phikon | TCGA-RCC | 0.964 (0.936-0.987) | 0.960 (0.937-0.982) | 0.995 (0.989-0.999) |
| Ctranspath | TCGA-RCC | 0.883 (0.820-0.938) | 0.888 (0.833-0.939) | 0.981 (0.965-0.992) |
| UNI | TCGA-RCC | 0.903 (0.847-0.954) | 0.913 (0.862-0.957) | 0.982 (0.966-0.995) |
| CONCH | TCGA-RCC | 0.941 (0.891-0.979) | 0.937 (0.896-0.974) | 0.988 (0.977-0.997) |
| PLIP | TCGA-RCC | 0.899 (0.845-0.947) | 0.904 (0.853-0.948) | 0.980 (0.963-0.993) |
| CHIEF | TCGA-RCC | 0.889 (0.824-0.945) | 0.900 (0.843-0.945) | 0.981 (0.964-0.993) |
| Prov-Gigapath | TCGA-RCC | 0.903 (838-0.958) | 0.907 (0.850-0.955) | 0.986 (0.975-0.996) |
| GPFM | TCGA-RCC | 0.925 (0.874-0.967) | 0.930 (0.885-0.966) | 0.996 (0.992-0.999) |
| ResNet50 | Center-3-RCC* | 0.490 (0.415-0.562) | 0.409 (0.332-0.496) | 0.747 (0.662-0.824) |
| Phikon | Center-3-RCC* | 0.433 (0.378-0.495) | 0.337 (0.254-0.422) | 0.831 (0.744-0.909) |
| Ctranspath | Center-3-RCC* | 0.735 (0.641-0.817) | 0.735 (0.640-0.821) | 0.910 (0.853-0.953) |
| UNI | Center-3-RCC* | 0.713 (0.614-0.799) | 0.717 (0.614-0.808) | 0.921 (0.868-0.959) |
| CONCH | Center-3-RCC* | 0.560 (0.472-0.647) | 0.521 (0.407-0.618) | 0.822 (0.745-0.890) |
| PLIP | Center-3-RCC* | 0.523 (0.453-0.593) | 0.451 (0.365-0.530) | 0.818 (0.743-0.883) |
| CHIEF | Center-3-RCC* | 0.757 (0.671-0.841) | 0.755 (0.658-0.840) | 0.921 (0.869-0.960) |
| Prov-Gigapath | Center-3-RCC* | 0.628 (0.532-0.718) | 0.607 (0.503-0.704) | 0.751 (0.678-0.817) |
| GPFM | Center-3-RCC* | 0.759 (0.667-0.835) | 0.756 (0.666-0.843) | 0.922 (0.868-0.963) |
@@
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | TCGA-BRCA | 0.658 (0.585-0.735) | 0.691 (0.596-0.783) | 0.846 (0.768-0.911) |
| Phikon | TCGA-BRCA | 0.794 (0.718-0.865) | 0.835 (0.751-0.901) | 0.936 (0.887-0.977) |
| Ctranspath | TCGA-BRCA | 0.843 (0.767-0.914) | 0.859 (0.790-0.917) | 0.931 (0.870-0.975) |
| UNI | TCGA-BRCA | 0.869 (0.797-0.929) | 0.879 (0.810-0.932) | 0.946 (0.894-0.987) |
| CONCH | TCGA-BRCA | 0.835 (0.750-0.905) | 0.875 (0.807-0.934) | 0.944 (0.902-0.979) |
| PLIP | TCGA-BRCA | 0.823 (0.747-0.897) | 0.820 (0.750-0.888) | 0.893 (0.824-0.950) |
| CHIEF | TCGA-BRCA | 0.790 (0.717-0.866) | 0.829 (0.756-0.896) | 0.899 (0.823-0.960) |
| Prov-Gigapath | TCGA-BRCA | 0.884 (0.821-0.940) | 0.877 (0.813-0.931) | 0.942 (0.884-0.983) |
| GPFM | TCGA-BRCA | 0.881 (0.813-0.947) | 0.907 (0.850-0.956) | 0.950 (0.898-0.990) |
| ResNet50 | Center-3-LD* | 0.441 (0.394-0.480) | 0.441 (0.406-0.472) | 0.562 (0.387-0.737) |
| Phikon | Center-3-LD* | 0.824 (0.613-1.000) | 0.849 (0.597-1.000) | 0.915 (0.726-1.000) |
| Ctranspath | Center-3-LD* | 0.814 (0.606-0.991) | 0.814 (0.578-0.957) | 0.814 (0.443-1.000) |
| UNI | Center-3-LD* | 0.750 (0.500-1.000) | 0.819 (0.491-1.000) | 0.874 (0.616-1.000) |
| CONCH | Center-3-LD* | 0.500 (0.500-0.500) | 0.472 (0.447-0.491) | 0.500 (0.500-0.500) |
| PLIP | Center-3-LD* | 0.637 (0.412-0.827) | 0.502 (0.374-0.635) | 0.614 (0.321-0.893) |
| CHIEF | Center-3-LD* | 0.500 (0.500-0.500) | 0.472 (0.447-0.491) | 0.811 (0.451-1.000) |
| Prov-Gigapath | Center-3-LD* | 0.652 (0.404-0.898) | 0.472 (0.441-0.736) | 0.784 (0.542-1.000) |
| GPFM | Center-3-LD* | 0.887 (0.686-0.991) | 0.837 (0.648-0.966) | 0.905 (0.694-1.000) |
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | BRACS | 0.568 (0.515-0.615) | 0.522 (0.463-0.571) | 0.835 (0.776-0.892) |
| Phikon | BRACS | 0.707 (0.621-0.797) | 0.701 (0.602-0.800) | 0.898 (0.852-0.942) |
| Ctranspath | BRACS | 0.674 (0.592-0.757) | 0.664 (0.559-0.754) | 0.908 (0.871-0.946) |
| UNI | BRACS | 0.746 (0.660-0.840) | 0.738 (0.640-0.824) | 0.913 (0.865-0.956) |
| CONCH | BRACS | 0.677 (0.606-0.752) | 0.668 (0.575-0.771) | 0.923 (0.883-0.958) |
| PLIP | BRACS | 0.679 (0.596-0.773) | 0.676 (0.579-0.782) | 0.866 (0.805-0.917) |
| CHIEF | BRACS | 0.717 (0.635-0.804) | 0.723 (0.617-0.814) | 0.922 (0.879-0.958) |
| Prov-Gigapath | BRACS | 0.720 (0.620-0.810) | 0.715 (0.613-0.802) | 0.927 (0.885-0.961) |
| GPFM | BRACS | 0.749 (0.660-0.834) | 0.758 (0.658-0.841) | 0.936 (0.896-0.965) |
| ResNet50 | Center-3-BRCA* | 0.618 (0.563-0.669) | 0.521 (0.472-0.564) | 0.624 (0.555-0.688) |
| Phikon | Center-3-BRCA* | 0.636 (0.584-0.691) | 0.539 (0.493-0.589) | 0.659 (0.607-0.713) |
| Ctranspath | Center-3-BRCA* | 0.583 (0.521-0.645) | 0.543 (0.493-0.596) | 0.670 (0.617-0.720) |
| UNI | Center-3-BRCA* | 0.556 (0.493-0.616) | 0.504 (0.454-0.552) | 0.614 (0.559-0.666) |
| CONCH | Center-3-BRCA* | 0.479 (0.450-0.503) | 0.178 (0.146-0.208) | 0.500 (0.498-0.501) |
| PLIP | Center-3-BRCA* | 0.578 (0.522-0.634) | 0.566 (0.514-0.618) | 0.681 (0.627-0.736) |
| CHIEF | Center-3-BRCA* | 0.568 (0.506-0.624) | 0.546 (0.495-0.597) | 0.670 (0.615-0.723) |
| Prov-Gigapath | Center-3-BRCA* | 0.587 (0.530-0.642) | 0.574 (0.521-0.629) | 0.678 (0.625-0.732) |
| GPFM | Center-3-BRCA* | 0.675 (0.625-0.720) | 0.543 (0.490-0.594) | 0.694 (0.641-0.747) |